OpenStudy (savannatigris):

Simplify using double angle formulas: 2 sin 35° cos 35°.

terenzreignz (terenzreignz):

You'd do well to recall that \[\Large \sin(2\alpha) = 2\sin(\alpha)\cos(\alpha)\]

OpenStudy (aonz):

btw the double angle formula is in a ratio is 2 to 1. just remember that :) sin4x = 2 sin2x cos2x